11/1/2019 The Piano Co. borrowed funds to replace its damaged roof
Amount borrowed              180,000
Interest rate 5%
Term of note 6 months
INSTRUCTIONS: Prepare the journal entries to record
the issuance of the note
the accrual of interest at year end
the payment of the note on its due date
On January 1, 2019, Teasdale Corp. issued 10 year              800,000 5% bond payable.
Interest is payable semi-annually on June 30 and December 31.
INSTRUCTIONS: Prepare the journal entry to record the issuance of these bonds assuming they
were issued at 98.
On December 31, 2019, Rockwell Inc. issues a long term note payable for              400,000
The note is a 10 year note with an interest rate of 8%.
The terms of the note call for semi-annual installment payments of                49,317
INSTRUCTIONS: Prepare the journal entries to record the note on 12-31-17 and to pay the
first installment payment on 6-30-2018.
Prepare the journal entry for 12-31-18 (second payment)
Answer #1
1 Date Account titles and explanation Debit Credit
11/1/2019. Cash 180000
Note payable 180000
(Issuance of the note)
12/31/2019 Interest expense (180000*5%*2/12) 1500
Interest payable 1500
(Accrual of interest at the year end for Nov and Dec
For 2 months)
04/30/2020. Note payable 180000
Interest expense (180000*5%*4/12) 3000
Interest payable 1500
Cash 184500
(Payment of the note on it's due date)
2 Issue price=Number of bonds*Issue price
Let's assume par value=$ 100
Number of bonds=Total face value/Par value=800000/100=8000
Issue price=Number of bonds*Issue price=8000*98=$ 784000
Discount on issue of bonds=Total face value-Issue price=800000-784000=$ 16000
Date Account titles and explanation Debit Credit
01/01/2019. Cash 784000
Discount on issue of bonds 16000
Bonds payable 800000
(Issuance of the bonds)
3 Date Account titles and explanation Debit Credit
12/31/2017. Cash 400000
Note payable 400000
(Issuance of the note)
06/30/2018. Interest expense (400000*8%*6/12) 16000
Note payable (49317-16000) 33317
Cash 49317
(First installment payment)
12/31/2018. Interest expense (400000-33317)*8%*6/12 14667
Note payable (49317-14667) 34650
Cash 49317
(Second installment payment)